# POINTS: Improving Your Vision-language Model with Affordable Strategies

### v1

@@ 4.4 Ablation Study and Analysis
As shown in Table 1, each component introduced in subsection 3.1 contributes to steady improvements. These enhancements are significant; for instance, after introducing Dynamic High Resolution to split the input image, we observe substantial improvements in OCR-related tasks, such as OCRBench, with performance increasing from 56.9% to 60.3%. Additionally, the use of high-resolution images with Dynamic High Resolution helps reduce hallucination, primarily due to the increased detail in the high-resolution images. Furthermore, replacing the original Dynamic High Resolution with CATTY results in notable improvements across various benchmarks, with OCR-related benchmarks showing greater gains than others. This is likely because image distortion has a more pronounced negative impact on text within images. Compared to general visual feature extraction, the ability to extract text features from images is limited for CLIP-ViT(Radford et al., 2021), as it was trained on a large quantity of general image-text pairs. Consequently, we observe substantial improvements on OCRBench after integrating features from an additional ViT, post-trained on text-rich images. Among the 5 strategies, incorporating more visual instruction tuning datasets by Individual Select yields the most significant improvements. This observation aligns with existing works(Chen et al., 2024c; Li et al., 2024; Tong et al., 2024), underscoring the importance of selecting effective datasets during the visual instruction tuning stage.
@@
GPT-4o-0513 - 61.3 55.0 736 84.6 69.1 63.9 69.2 90.7 2310.3 75.4 102.0
Claude3.5-Sonnet - 61.6 49.9 788 80.2 66.0 62.2 65.9 88.9 1920.0 60.1 81.0
Gemini-1.5-Pro - 57.7 45.6 754 79.1 64.0 59.1 60.6 85.7 2110.6 64.1 95.3
@@
Cambrian-34B 81.4 50.3 41.6 591 79.5 53.2 54.2 50.4 85.6 2049.9 67.1 82.0
Ovis1.5-LLaMA3-8B - 63.0 45.0 744 82.5 50.9 57.3 48.3 88.8 1948.5 64.2 79.9
InternVL2-8B - 58.3 45.0 794 83.6 54.3 61.5 51.2 97.1 2215.1 64.2 73.3
IXC-2.5 - 63.7 43.1 686 81.6 49.3 59.9 42.9 96.6 2233.1 67.8 70.2
@@
Table 5: Comparison between different methods. CF: CapFusion(Yu et al., 2024), DHR: Dynamic high resolution(Chen et al., 2024c), CATTY: Consistent aspect ratio dynamic high resolution proposed by us, DEV: Dual vision encoder(Wei et al., 2023), IS: Individual select(Liu et al., 2024c). MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), SCI: ScienceQA(Lu et al., 2022), MME: MME(Yin et al., 2023), RWQ: RealWorldQA, Wild: LLaVA-Wild(Liu et al., 2024b). Cambrian-34: Cambrian-34B(Tong et al., 2024), Ovis1.5-LLaMA3-8B: Ovis1.5(Lu et al., 2024b), IXC-2.5: IXC-2.5(Zhang et al., 2024), OneVision: LLaVA-OneVision(Li et al., 2024). The language model POINTS-9B uses is Yi-1.5-9B (Young et al., 2024).
@@
As described in previous sections, increasing adding more instruction tuning datasets often reaches a plateau where further increasing the number of datasets yields minimal improvement. However, by incorporating model soup across different datasets, we observe substantial enhancements, as shown in Table 5, with the overall score increasing from 59.0 to 61.2. We also compare the benefits of various model soup strategies. Among them, greedy soup achieves the best performance, outperforming maximum soup and average soup by 0.6 and 0.4 points, respectively. Unless otherwise specified, we will use greedy soup by default in subsequent experiments. Additionally, we include the results of conducting model soup over different hyperparameters, e.g. different learning rates. As shown, model soup over hyperparameters brings only marginal improvement; detailed results are provided in the appendix. Furthermore, we verify in the appendix that model soup consistently improves performance regardless of the Base Set used.

### v2

@@ 4.4 Ablation Study and Analysis
As shown in Table 1, each component introduced in subsection 3.1 contributes to steady improvements. These enhancements are significant; for instance, after introducing Dynamic High Resolution to split the input image, we observe substantial improvements in OCR-related tasks, such as OCRBench, with performance increasing from 49.6% to 55.6%. Additionally, the use of high-resolution images with Dynamic High Resolution helps reduce hallucination, primarily due to the increased detail in the high-resolution images. Furthermore, replacing the original Dynamic High Resolution with CATTY results in notable improvements across various benchmarks, with OCR-related benchmarks showing greater gains than others. This is likely because image distortion has a more pronounced negative impact on text within images. Compared to general visual feature extraction, the ability to extract text features from images is limited for CLIP-ViT(Radford et al., 2021), as it was trained on a large quantity of general image-text pairs. Consequently, we observe substantial improvements on OCRBench after integrating features from an additional ViT, post-trained on text-rich images. Among the 5 strategies, incorporating more visual instruction tuning datasets by Individual Select yields the most significant improvements. This observation aligns with existing works(Chen et al., 2024c; Li et al., 2024; Tong et al., 2024), underscoring the importance of selecting effective datasets during the visual instruction tuning stage.
@@
GPT-4o-0513 - 61.3 55.0 73.6 84.6 69.1 63.9 69.2 90.7 2310.3 75.4 102.0
Claude3.5-Sonnet - 61.6 49.9 78.8 80.2 66.0 62.2 65.9 88.9 1920.0 60.1 81.0
Gemini-1.5-Pro - 57.7 45.6 75.4 79.1 64.0 59.1 60.6 85.7 2110.6 64.1 95.3
@@
Cambrian-34B 81.4 50.3 41.6 59.1 79.5 53.2 54.2 50.4 85.6 2049.9 67.1 82.0
Ovis1.5-LLaMA3-8B - 63.0 45.0 74.4 82.5 50.9 57.3 48.3 88.8 1948.5 64.2 79.9
InternVL2-8B - 58.3 45.0 79.4 83.6 54.3 61.5 51.2 97.1 2215.1 64.2 73.3
IXC-2.5 - 63.7 43.1 68.6 81.6 49.3 59.9 42.9 96.6 2233.1 67.8 70.2
@@
Table 5: Comparison between different methods. MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), SCI: ScienceQA(Lu et al., 2022), MME: MME(Yin et al., 2023), RWQ: RealWorldQA, Wild: LLaVA-Wild(Liu et al., 2024b). Cambrian-34: Cambrian-34B(Tong et al., 2024), Ovis1.5-LLaMA3-8B: Ovis1.5(Lu et al., 2024b), IXC-2.5: InternLM-XComposer-2.5(Zhang et al., 2024), OneVision: LLaVA-OneVision(Li et al., 2024). The language model POINTS-9B uses is Yi-1.5-9B (Young et al., 2024).
@@
As described in previous sections, increasing adding more instruction tuning datasets often reaches a plateau where further increasing the number of datasets yields minimal improvement. However, by incorporating model soup across different datasets, we observe substantial enhancements, as shown in Table 5, with the overall score increasing from 59.0 to 61.8. We also compare the benefits of various model soup strategies. Among them, greedy soup achieves the best performance, outperforming maximum soup and average soup by 0.8 and 0.6 points, respectively. Unless otherwise specified, we will use greedy soup by default in subsequent experiments. Additionally, we include the results of conducting model soup over different hyperparameters, e.g. different learning rates. As shown, model soup over hyperparameters brings only marginal improvement; detailed results are provided in the appendix. Furthermore, we verify in the appendix that model soup consistently improves performance regardless of the Base Set used.